# Mathematical Politics

### v1

@@ Abstract
Politics today is largely about the art of messaging to influence the public, but the mathematical theory of messaging — information and communication theory — can turn this art into a precise analysis, both qualitative and quantitative, that enables us to gain retrospective understandings of political events and to make forward-looking predictions.
@@ Introduction
[…]
@@
What these numbers tell us is that if an election were to take place tomorrow, then the probability of candidate $A$ winning the election is about 100%. But they say very little about what happens in a future election. If the election were to take place in a year, then most people will think that the likelihood of candidate $A$ winning the election is close to 52%. But what if the election is to take place in six months, or in three months? To interpolate today’s statistics with future statistics, we need a dynamical model. In the present context, what generates the dynamics is the revelation of information, so let me take the simplest nontrivial model $\xi_{t}=\sigma Xt+\epsilon_{t}$ for the flow of information. Then we can generate a model for the dynamics of the poll statistics $\pi_{it}={\mathbb{P}}(X=x_{i}|\xi_{\leq t})$, where $\xi_{\leq t}$ denotes information gathered up to time $t$. Using the Bayes formula we find, in this example,
@@
for the a posteriori support rates. This is how we can use a model for the flow of information to interpolate between today and the future. If the election is to take place at a future date $T$, then the vote share of candidate $j$ exceeding a threshold value $K$ on the election day, based on the knowledge today and on how information flows to the electorates from today till the election day, can be determined by working out ${\mathbb{P}}(\pi_{jT}>K)$. As an example, in the case of a two-candidate race (Brody 2019), writing $p_{1}=p$ and $p_{2}=1-p$, a calculation shows that
@@
[…]
@@
[…]
Here, $X_{t}(\psi)=\sum_{i}x_{i}(\psi^{i}_{t})^{2}/\sum_{i}(\psi^{i}_{t})^{2}$ is the conditional mean and $V_{t}(\psi)=\sum_{i}(x_{i}-X_{t})^{2}(\psi^{i}_{t})^{2}/\sum_{i}(\psi^{i}_{t})^{2}$ is the conditional variance of the random variable $X$; both are functions on ${\mathfrak{M}}$. The drift vector field of $\psi^{i}_{t}$, when there are three candidates, is plotted in figure 2.
@@
[…]
@@
Figure 3: If $p_{1}=0.01$ and $p_{2}=0.99$, then even if the noisy information is provided by $\xi_{t}=\sigma x_{1}t+\epsilon_{t}$, indicating that $X=x_{1}$ is the correct alternative so that $\pi_{1t}$ should approach $1$, a Bayesian thinker will maintain the position $\pi_{1t}\approx 0$ for a long time. Four sample paths are shown here for $\sigma=1$.
@@
where $\rho$ is the correlation of the two Brownian noises $\epsilon^{1}_{t}$ and $\epsilon^{2}_{t}$, and $\sigma$ is a certain positive function of the parameters $\sigma_{1}$, $\sigma_{2}$, and $\rho$. We can think of $\xi^{1}_{t}$ and $\xi^{2}_{t}$ representing, for instance, information emerging from two competing political parties. Suppose that $\rho>0$ and that $\xi_{t}^{2}$, say, contains disinformation. In this case, by letting $\sigma_{1}>\sigma_{2}/\rho$, it is possible to reverse the impact of disinformation, that is, fake news will turn against its spreader. The lesson for the first political party therefore is to deliver clear and transparent messages to enhance $\sigma_{1}$, rather than to pay too much attention to counter the disinformation.
@@
In cognitive psychology, for long it was taken for granted, quite understandably, that all propositions are compatible. One consequence of this assumption is that when people’s behaviour deviates away from what is predicted by the rules of probability, for example a violation of the Bayesian updating, such a behaviour can be classified as irrational. A classic example is the notorious conjunction fallacy (Tversky & Kahneman 1983) , whereby “$A$ and $B$” is thought more likely than“$A$”. According to standard logic and probability theory this is never true, but in experiment after experiment it can be human-true. But standard logic and probability theory is applicable only so far as when $A$ and $B$ admit a joint probability law.
It is legitimate therefore to ask: What if $A$ and $B$ were not compatible? This is in effect the question addressed by Jerome Busemeyer and others in cognitive science (see Busemeyer & Bruza 2024, and references cited therein). The answer they have found was remarkable. Much of the seemingly irrational human behaviours reported in the cognitive psychology literature, for which there are no straightforward explanations using only compatible propositions, could be explained consistently using models accommodating incompatible propositions. Statistical analysis of random events involving incompatible propositions, however, is what quantum theory is about. Indeed, there are two distinctions between Kolmogorovean classical probability theory and what is known as quantum probability theory. The first is that the square-root probabilities $\psi_{i}$, called probability amplitudes in quantum mechanics, can take complex values so that the probabilities are recovered using modulus squares: $p_{i}=|\psi_{i}|^{2}$. (Imaginary number plays two roles in quantum mechanics, namely, to rotate plane waves by the right angle, and to set the direction of time. It is known that quantum theory cannot be formulated without the use of complex numbers, whereas it is not known whether complex numbers are required for a consistent characterisation of cognitive behaviours.) The second is that random variables, which are modelled on Hilbert space as matrices, are in general not compatible. With the realisation that quantum probability models can more accurately and succinctly model cognitive behaviours, a new field emerged within cognitive science called ‘quantum cognition’, which is now an active field of research.
@@
The relevance of what I call here noncommutative communication theory in politics has long been appreciated by seasoned politicians and political advisors, purely on the empirical or intuitive grounds, as an art of messaging. Perhaps a simple anecdotal example suffices to highlight this. Take the 2016 ‘Brexit’ referendum in the United Kingdom on membership of the European Union. Some ‘leave’ campaigners put to the public the question, before the referendum, whether they want more funding for the National Health Service, with the implication, true or false, that leaving the European Union implies more funding for the health service. If the membership question and the health service funding question were compatible, then classical probability theory shows that asking the funding question will have no implication in the referendum outcome. Those leave campaigners believed otherwise, and the result was ‘out’.
There are innumerable such examples whereby political campaigners instinctively know how to shift the frame of reference of the public opinion, but the science underpinning this effect is only emerging in the form of quantum cognition and noncommutative communication theory, over the past two decades. Indeed, even in the hypothetical scenario whereby a person believes 100% in an alternative, there is a way of changing the view of that person – an impossibility in classical probability. The trick is not to convince the person otherwise, which is futile because a counterargument to that alternative is compatible with the argument itself. Counterintuitive this may be, a counterargument will have no effect on the view of that person. Instead, by asking a single incompatible question, and importantly a single question is sufficient here, it is possible to change the belief of the person on the original issue to a number strictly less than 100%.
@@ Discussion
There are new mathematical challenges that emerge as well. For example, in traditional communication theory, there are no ambiguities as for the interpretation of the signal (the random variable $X$ in the current example). In politics, however, there is a huge ambiguity. Consider a newspaper article on an issue. The view of a person on this issue may be modelled by $X$ (taking the values $\{x_{i}\}$ with the probabilities $\{p_{i}\}$) so that the information contained in the article can be modelled, say, by $\xi=\sigma X+\epsilon$; but only for that person. Another reader of the same article will have a different view on the issue, represented by $Y$ (also taking the values $\{x_{i}\}$ but with the probabilities $\{q_{i}\}$), so the information model for that person takes the form, say, $\xi^{\prime}=\sigma Y+\epsilon$, even though we are speaking about an identical news article. Thus, information content of a message, that is, the amount of information contained in the message about the value of the signal, known in communication theory as mutual information (Gelfand & Yaglom 1957), is receiver dependent. In a time series version of the problem, statistical analysis requires the specification of information in the form of conditioning, but when there are ambiguities regarding which, or whose information to consider, it is not clear how to conduct such analysis, and this is a challenge in pure mathematics.

### v2

@@ Abstract
Politics today is largely about the art of messaging to influence the public, but the mathematical theory of messaging — information and communication theory — can turn this art into a precise analysis, both qualitative and quantitative, that enables us to gain retrospective understandings of past political events and to make forward-looking future predictions.
@@ Introduction
In communication theory there is a standard procedure to isolate the arrival of new information from an information-providing time series $\{\xi_{t}\}$. To understand this, consider the information gathered over the time interval $t,t+\mbox{$\rm d$}t]$, which is represented by the increment $\mbox{$\rm d$}\xi_{t}$. If we write ${\mathbb{E}}_{t}[-]$ for the expectation conditional on the information available up to time $t$, then the predicted arrival of new information at time $t$ is given by ${\mathbb{E}}_{t}[\mbox{$\rm d$}\xi_{t}]$. It then follows that the discrepancy between the predicted and realised increments
$\mbox{$\rm d$}W_{t}=\mbox{$\rm d$}\xi_{t}-{\mathbb{E}}_{t}[\mbox{$\rm d$}\xi_{t}]$
@@
What these numbers tell us is that if an election were to take place tomorrow, then the probability of candidate $A$ winning the election is about 100%. But they say very little about what happens in a future election. If the election were to take place in a year, then most people will think that the likelihood of candidate $A$ winning the election is close to 52%. But what if the election is to take place in six months, or in three months? To interpolate today’s statistics with future statistics, we need a dynamical model. In the present context, what generates the dynamics is the revelation of information, so let me take the simplest nontrivial model $\xi_{t}=\sigma Xt+\epsilon_{t}$ for the flow of information. Then we can generate a model for the dynamics of the poll statistics by the conditional probability $\pi_{it}={\mathbb{P}}(X=x_{i}|\xi_{\leq t})$, where $\xi_{\leq t}$ denotes information gathered up to time $t$. Using the Bayes formula we find, in this example,
@@
for the a posteriori support rates. This is how we can use a model for the flow of information to interpolate between today and the future. If the election is to take place at a future date $T$, then the vote share of candidate $j$ exceeding a threshold value $K$ on the election day, based on the knowledge today and on how information flows to the electorates from today till the election day, can be determined by working out the probability ${\mathbb{P}}(\pi_{jT}>K)$. As an example, in the case of a two-candidate race (Brody 2019), writing $p_{1}=p$ and $p_{2}=1-p$, a calculation shows that
@@
Let me now discuss an important dynamical characteristic of the Bayesian updating leading to the transformation $p_{i}\to\pi_{it}$ from the prior to the posterior. For this purpose it will be useful to take the derivative of $\pi_{it}$, making use of the Ito stochastic calculus. Specifically, $\pi_{it}$ is a function of two variables $t$ and $\xi_{t}$, so its (stochastic) derivative $\mbox{$\rm d$}\pi_{it}$ can be worked out by expanding it as a Taylor series and retain terms of first order, bearing in mind that $(\mbox{$\rm d$}\xi_{t})^{2}=\mbox{$\rm d$}t$. If we then substitute $\mbox{$\rm d$}\xi_{t}=\mbox{$\rm d$}W_{t}+{\mathbb{E}}_{t}[\mbox{$\rm d$}\xi_{t}]$ into the result, we find that $\mbox{$\rm d$}\pi_{it}\propto\mbox{$\rm d$}W_{t}$. That is, increments of the posterior probability are proportional to those of the innovation process. Such a process is known as a ‘martingale’, representing a conserved quantity, albeit on average.
@@
$\mbox{$\rm d$}\psi^{i}_{t}=-\textstyle{\frac{1}{16}}\nabla^{i}V_{t}\,\mbox{$\rm d$}t+\textstyle{\frac{1}{4}}\nabla^{i}X_{t}\,\mbox{$\rm d$}W_{t}.$
Here, $X_{t}(\psi)$ is the conditional mean and $V_{t}(\psi)$ is the conditional variance of the random variable $X$; both are functions on $\mathfrak{M}$}. The drift vector field of the state $\psi^{i}_{t}$, which determines the orientational direction, when there are three candidates, is plotted in figure 2.
@@
An Ito process $\{x_{t}\}$ is defined by the equation $\mbox{$\rm d$}x_{t}=\mu_{t}\mbox{$\rm d$}t+\sigma_{t}\mbox{$\rm d$}W_{t}$, where $\{\mu_{t}\}$ and $\{\sigma_{t}\}$ are called the drift and the volatility of $\{x_{t}\}$. We can consider an Ito process $\{x_{t}\}$ taking values on a manifold $\mathfrak{M}$}, driven by an $m$-dimensional Wiener process $\{W_{t}^{i}\}_{i=1,\ldots,m}$. Let $\nabla_{a}$ be a torsion-free connection on $\mathfrak{M}$} such that for any vector field $\zeta^{a}$ its covariant derivative in local coordinates is $\delta_{\bf b}^{b}\delta_{a}^{\bf a}(\nabla_{b}\zeta^{a})=\frac{\partial\zeta^{\bf a}}{\partial x^{\bf b}}+\Gamma_{\bf bc}^{\bf a}\zeta^{\bf c},$ where $\delta_{a}^{\bf a}$ is the standard coordinate basis in a given coordinate patch. Suppose we have an Ito process taking values in $\mathfrak{M}$}. Let $x_{t}^{\bf a}$ denote the coordinates of the process in a particular patch. Then writing $h^{\bf ab}=\sigma_{i}^{\bf a}\sigma^{{\bf b}i}$ we define the drift process $\mu^{\bf a}$ by $\mu^{\bf a}\mbox{$\rm d$}t=\mbox{$\rm d$}x^{\bf a}+\mbox{$\textstyle\frac{1}{2}$}\Gamma_{\bf bc}^{\bf a}h^{\bf bc}\mbox{$\rm d$}t-\sigma_{i}^{\bf a}\mbox{$\rm d$}W_{t}^{i}$. Alternatively, we can write the covariant Ito differential as $\mbox{$\rm d$}x^{a}=\delta_{\bf a}^{a}(\mbox{$\rm d$}x^{\bf a}+\mbox{$\textstyle\frac{1}{2}$}\Gamma_{\bf bc}^{\bf a}h^{\bf bc}\mbox{$\rm d$}t)$, where $\delta^{a}_{\bf a}$ is the dual coordinate basis. Then we have $\mbox{$\rm d$}x^{a}=\mu^{a}\mbox{$\rm d$}t+\sigma^{a}_{i}\mbox{$\rm d$}W^{i}_{t}$ for the covariant Ito differential on $\mathfrak{M}$} associated with the given connection.
@@
Figure 3: If $p_{1}=0.01$ and $p_{2}=0.99$, then even if the noisy information is provided by $\xi_{t}=\sigma x_{1}t+\epsilon_{t}$, indicating that $X=x_{1}$ is the correct alternative so that $\pi_{1t}$ should approach 11, a Bayesian thinker will maintain the position $\pi_{1t}\approx 0$ for a long time. Four sample paths are shown here for $\sigma=1$.
@@
where $\rho$ is the correlation of the two Brownian noises $\epsilon^{1}_{t}$ and $\epsilon^{2}_{t}$, and $\sigma$ is a certain positive function of the parameters $\sigma_{1}$, $\sigma_{2}$, and $\rho$. We can think of $\xi^{1}_{t}$ and $\xi^{2}_{t}$ representing, for instance, information emerging from two competing political parties. Suppose that $\rho>0$ and that $\xi_{t}^{2}$, say, contains disinformation. In this case, by letting $\sigma_{1}>\sigma_{2}/\rho$, it is possible to reverse the impact of disinformation, that is, fake news will turn against its spreader (Brody & Yuasa 2025). The lesson for the first political party therefore is to deliver clear and transparent messages to enhance $\sigma_{1}$, rather than to pay too much attention to counter the disinformation.
@@
Transmission, feedback, and control
Thus far our attention has been focused on how noisy information is processed and analysed, which we used to deduce the resulting action. Given the fact that this analysis generates an uncertainty-minimising flow, it is natural to ask whether an individual can transmit information back into the system in the form of a feedback, with the view to further reduce uncertainties. That is, we can think of the possibility of feedback control. To this end consider an individual (the system), the environment (community, society, or a section of the society), and the communication channel (Internet, Twitter / X, Instagram, TikTok, and so on).
It is a well-known fact that channel capacity cannot be increased by feedback. Nevertheless, it turns out that noise-free feedback from the system to the environment can enhance the information gained from the environment. Consider the following, somewhat idealised setup of Davis (1978). In this setup, the information entering the system from the environment is obscured by noise, but the environment can detect the signal transmitted from the system without noise. The system (individual) is interested in the amount of information contained in the message coming from the environment about the value of the signal. This quantifier is known in communication theory as mutual information (Gelfand & Yaglom 1957), which is determined by the likelihood ratio for detecting the signal. The question therefore is what feedback should the individual transmit back into the community so as to maximise the mutual information. In the case of a Gaussian signal, the optimal transmission turns out to be the increments $\mbox{$\rm d$}W_{t}$ of the innovation process itself (Davis 1978). That is, only new information is transmitted from the system to the environment to maximally reduce the uncertainty of the signal.
There is a profound implication of this result, albeit in a slightly idealised setup, in today’s digital society. From individual’s perspective, transmission of information is time and energy consuming. Hence in practice unless the magnitude $|\mbox{$\rm d$}W_{t}|$ of the new information is above a threshold, it is unlikely that this is fed back into the community. Recall that the new information $\mbox{$\rm d$}W_{t}$ is the discrepancy $\mbox{$\rm d$}\xi_{t}-{\mathbb{E}}_{t}[\mbox{$\rm d$}\xi_{t}]$ between the predicted and realised increments of the information. Now suppose that the information $\mbox{$\rm d$}\xi_{t}$ over the time interval $t,t+\mbox{$\rm d$}t]$ is secretly contaminated with disinformation $\Delta f$. Then this will shift the realised increment $\mbox{$\rm d$}\xi_{t}$ by $\Delta f$ but not the predicted increment ${\mathbb{E}}_{t}[\mbox{$\rm d$}\xi_{t}]$, because the receiver is unaware of the existence of disinformation. It follows that for sufficiently large $\Delta f$, the likelihood of the distorted innovation $|\mbox{$\rm d$}W_{t}+\Delta f|$ exceeding the threshold value is very high.
It is known empirically that the likelihood of a person disseminating disinformation is significantly higher than that of the person disseminating true information (Vosoughi et al. 2018). The observation here demonstrates the intuitively acknowledged fact that the higher the novelty of information, the more likely it is that the information is transmitted. Importantly, though, this is the consequence of a highly rational behaviour, following the rule of optimal feedback control. The consequence therefore is to further enhance noise in the system, and this enhancement is amplified in a society that is becoming more reliant on the use of artificial intelligence systems.
@@
In cognitive psychology, for long it was taken for granted, quite understandably, that all propositions are compatible. One consequence of this assumption is that when people’s behaviour deviates away from what is predicted by the rules of probability, for example a violation of the Bayesian updating, such a behaviour can be classified as irrational. A classic example is the notorious conjunction fallacy (Tversky & Kahneman 1983) , whereby “$A$ and BB” is thought more likely than“AA”. According to standard logic and probability theory this is never true, but in experiment after experiment it can be human-true. But standard logic and probability theory is applicable only so far as when $A$ and $B$ admit a joint probability law.
It is legitimate therefore to ask: What if $A$ and $B$ were not compatible? This is in effect the question addressed by Jerome Busemeyer and others in cognitive science (see Busemeyer & Bruza 2024, and references cited therein). The answer they have found was remarkable. Much of the seemingly irrational human behaviours reported in the cognitive psychology literature, for which there are no straightforward explanations using only compatible propositions, could be explained consistently using models accommodating incompatible propositions. Statistical analysis of random events involving incompatible propositions, however, is what quantum theory is about. Indeed, there are two fundamental distinctions between Kolmogorovean classical probability theory and what is known as quantum probability theory. The first is that the square-root probabilities $\psi_{i}$, called probability amplitudes in quantum mechanics, can take complex values so that the probabilities are recovered using modulus squares: $p_{i}=|\psi_{i}|^{2}$. (Imaginary number plays two roles in quantum mechanics, namely, to rotate plane waves by the right angle, and to set the direction of time. It is known that quantum theory cannot be formulated without the use of complex numbers, whereas it is not known whether complex numbers are required for a consistent characterisation of cognitive behaviours.) The second is that random variables, which are modelled on Hilbert space as matrices, are in general not compatible. With the realisation that quantum probability models can more accurately and succinctly model cognitive behaviours, a new field emerged within cognitive science called ‘quantum cognition’, which is now an active field of research.
@@
The relevance of what I call here noncommutative communication theory in politics has long been appreciated by seasoned politicians and political advisors, purely on the empirical or intuitive grounds, as an art of messaging. Perhaps a simple anecdotal example suffices to highlight this. Take the 2016 ‘Brexit’ referendum in the United Kingdom on membership of the European Union. Some ‘leave’ campaigners put to the public the question, before the referendum, whether they want more funding for the National Health Service, with the implication, true or false, that leaving the European Union leads to more funding for the health service. If the membership question and the health service funding question were compatible, then classical probability theory shows that asking the funding question will have no implication in the referendum outcome. Those leave campaigners believed otherwise, and the result was ‘out’.
There are innumerable such examples whereby political campaigners instinctively know how to shift the frame of reference of the public opinion, but the science underpinning this effect is only emerging in the form of quantum cognition and noncommutative communication theory, over the past two decades. Indeed, even in the hypothetical scenario whereby a person believes 100% in an alternative, there is a way of changing the view of that person – an impossibility in classical probability (Basieva et al. 2017). The trick is not to convince the person otherwise, which is futile because a counterargument to that alternative is compatible with the argument itself. Counterintuitive this may be, a counterargument will have no effect on the view of that person. Instead, by asking a single incompatible question, and importantly a single question is sufficient here, it is possible to change the belief of the person on the original issue to a number strictly less than 100%.
@@ Discussion
There are new mathematical challenges that emerge as well. For example, in traditional communication theory, there are no ambiguities as for the interpretation of the signal (the random variable $X$ in the current example). In politics, however, there is a huge ambiguity. Consider a newspaper article on an issue. The view of a person on this issue may be modelled by $X$ (taking the values $\{x_{i}\}$ with the probabilities $\{p_{i}\}$) so that the information contained in the article can be modelled, say, by $\xi=\sigma X+\epsilon$; but only for that person. Another reader of the same article will have a different view on the issue, represented by $Y$ (also taking the values $\{x_{i}\}$ but with the probabilities $\{q_{i}\}$), so the information model for that person takes the form, say, $\xi^{\prime}=\sigma Y+\epsilon$, even though we are speaking about an identical news article. Thus, mutual information of a message, that is, the amount of information contained in the message about the value of the signal, is receiver dependent. In a time series version of the problem, statistical analysis requires the specification of information in the form of conditioning, but when there are ambiguities regarding which, or whose information to consider, it is not clear how to conduct such analysis, and this is a challenge already in pure mathematics.